Ray has some change in his pocket: four pennies, two nickels, four dimes,
two quarters, and two half dollars. He draws out a coin at random. The
probability of drawing any one of the coins is
0.06 for each penny.
0.01 for each dime.
0.03 for each nickel.
0.09 for each quarter.
0.24 for each half-dollar.
(a) Find the probability that the coin drawn is a dime.
(b) Find the probability that the coin drawn is a quarter or a half-dollar.
(c) Find the probability that the coin drawn is a nickel.
(d) Find the probability that the value of the coin drawn is less than 10 cents.

Answer by Edwin McCravy(20055)

Ray has some change in his pocket: four pennies, two nickels, four dimes,
two quarters, and two half dollars. He draws out a coin at random. The
probability of drawing any one of the coins is
0.06 for each penny.
So since he has four pennies, his probability
of drawing a penny is 4×0.06 = 0.24

0.01 for each dime.
So since he has four dimes, his probability
of drawing a dime is 4×0.01 = 0.04

0.03 for each nickel.
So since he has two nickels, his probability
of drawing a nickel is 2×0.03 = 0.06

0.09 for each quarter.
So since he has two quarters, his probability
of drawing a quarter is 2×0.09 = 0.18

0.24 for each half-dollar.
So since he has two half-dollars, his probability
of drawing a half-dollar is 2×0.24 = 0.48

Now, let's see if the problem is legitimate by
adding up all those probabilities and seeing if
we get 1.00:

0.24 = probability of drawing a penny.
0.04 = probability of drawing a dime.
0.06 = probability of drawing a nickel.
0.18 = probability of drawing a quarter.
0.48 = probability of drawing a half-dollar.
----
1.00

Yes, we get 1, the probability of certainty, since
it is certain that he will get a penny or a dime or
a nickel or a quarter or a half-dollar, so this is 
a legitimate probability problem.

(a) Find the probability that the coin drawn is a dime.
0.04

(b) Find the probability that the coin drawn is a quarter
or a half-dollar.
0.18 + 0.48 = 0.66 

(c) Find the probability that the coin drawn is a nickel.
0.06

(d) Find the probability that the value of the coin drawn is
less than 10 cents.
That's the same as if it had read this way: 

Find the probability that the coin drawn is a penny or a 
nickel.

0.24 + 0.06 = 0.30

Edwin